APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday August 15, 2010 (MJD 55424) ---=== OBSERVING TEAM ===--- Viktor M Steph Snedden Paul Harding Craig Loomis (via STUI messenger) ---=== OBSERVING PLAN ===--- Shakedown: Guider work, with possible MARVELS science later. ---=== OBSERVING SUMMARY ===--- High humidity, nearby thunderstorms prevented opening. Craig worked on guider software and we worked on guider tests inside the enclosure. We checked for shifting of the guide fiber centroids as a function of altitude and rotator position. Some shifting was seen, but tightening down the guide camera lens set screws in the barrel seemed to reduce the shift. Paul continues to analyze the data. ---=== OBSERVING LOG ===--- Afternoon: ---------- -----Status --------- 2010-08-16 00:20:11.537628 Guider cart=11 pl=4097 side=A (loaded) MCP instr 18 (mounted) motorPosition sp1: 18999 18998 19001 sp2: 7697 7800 7760 Median spectrographs temperatures: Tsp1=15.5 Tsp2=14.9 Outside air temperature=19.8C cwPositions=201,201,200,201 focus of secondary=0.0 Offset 0=PVT(0.0, 0.0, 4788618872.89) Offset 1=PVT(0.0, 0.0, 4788618872.89) gcameraTemp reported=-40.0 testActors, 2010-08-16 00:19:43 alerts ping OK alerts status OK apo ping OK apo status OK boss status OK gcamera ping OK gcamera status OK ecamera ping OK ecamera status OK guider ping OK guider status OK hub status OK mcp ping OK msg just testing OK perms status OK platedb ping OK platedb status OK sop ping OK sop status OK sos ping ** FAILED ** tcc show status OK tcc show time OK keys getFor=hub version OK -------- boss ping --??? check -------- 00:45Z - all system checked 00:48Z - opened doors, louvers, fans on and started ventillation. Night: ------ 01:45Z - Humidity climbing and storms are nearby. We stay closed with ventilation on. 01:50Z - The blip in the sigma on the cloud camera was a person, not clouds. 02:49Z - ThAr lamp on 03:39Z - Mounted cartridge 11 for guider tests - flexure tests and actor stability. Note: Ecamera actor segfaults when the ecamera is disconnected. Craig and Paul looked into that. 03:47Z - Craig started some simulator tests to see if the changes made to the gcamera actor that allows the ecamera to run on modified gcamera code has not broken the gcamera code. 05:12Z - Ventilation off. Humidity over the closure limit. We then started guide camera flexure tests: All guider exposure times are 0.5 s. The summary of exposures is below. The first sequence of exposures showed evidence of "flop" based on rotator position. We then went down and snugged down the screws on the guider's lens tube, under the assumption it was the lenses moving a bit and causing the shift in the guider flat images. The second set of exposures (after "snugging") showed an improvement, with reduced "flop". The data below are AFTER snugging down the lens set screws. The x and y positions below are those of the fiber centroid. gimg Alt Rot fiber 16 fiber 4 -00xx (mount) x y x y 16 0 0 616.86 889.74 454.30 191.72 17 0 0 616.86 889.73 454.29 191.71 19 60 60 616.42 890.62 453.84 192.59 20 60 60 616.45 890.66 453.83 192.63 21 60 120 616.87 889.97 454.24 192.02 22 60 180 616.57 890.69 453.97 192.77 23 60 240 616.48 890.95 453.93 192.96 24 60 300 616.53 891.33 453.95 193.34 25 60 360 616.29 891.29 453.68 193.26 26 60 300 616.62 890.92 454.03 192.94 27 60 240 616.50 891.04 453.93 193.07 28 60 180 616.47 891.08 453.85 193.14 29 60 120 616.54 891.21 453.87 193.27 30 60 60 616.42 891.34 453.76 193.32 31 60 0 616.43 891.24 453.83 193.19 32 60 0 616.43 891.24 453.83 193.20 33 60 -60 616.58 891.12 453.99 193.13 34 60 -120 616.46 891.13 453.90 193.16 35 60 -180 616.45 891.11 453.83 193.18 36 60 -180 616.44 891.15 453.82 193.19 37 60 0 616.20 891.67 453.57 193.64 38 30 0 616.30 891.33 453.69 193.30 39 30 0 616.28 891.29 453.66 193.27 40 30 60 616.54 891.23 453.88 193.20 41 30 120 616.88 891.34 454.20 193.39 42 30 180 616.57 892.25 453.93 194.26 43 30 240 616.29 892.77 453.68 194.75 44 30 300 616.36 892.47 453.73 194.47 45 30 360 616.27 892.14 453.62 194.09 46 30 240 616.33 892.37 453.73 194.39 47 30 120 616.45 892.73 453.74 194.77 48 30 0 616.37 892.38 453.73 194.34 49 30 -60 616.44 892.17 453.82 194.15 50 30 -120 616.33 892.49 453.72 194.51 51 30 -180 616.30 892.89 453.64 194.93 10:23Z - Nearby lightning and high humidity. We give up hope of opening. 10:25Z - ThAr lamp off endNight finished. ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:40Z 65 52 14 76 (ENE) -9999 - Partly cloudy No more weather entries - please check weather actor ---=== TELESCOPE STATUS ===--- 00:48Z Doors and louvers opened, fans on. 05:12Z Ventilation off 10:29ZZ Amplifiers off Altitude interock bypassed Status at 10:44Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cart 11 Counterweights at: 200 LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Sp_01 89 lbs, 22 psi Sp_02 62 lbs, 21 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_5 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: 61600 packets sent successfully getVersions, 2010-08-16 00:18:59 ----STUI software STUI: 1.0b2 2010-07-21 alerts: trunk+svn114742 apo: trunk+svn114701 boss: trunk+svn115542M boss.daq: 1.2.14 boss.specMech: 3.0.1 gcamera: trunk+svn115486 guider: branch_dstn-fiber-centers+svn115457M hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: trunk+svn111739 sop: v1_1_3 sos: ** FAILED ** tcc: 2.13.0 -----Other software sdssProcedure: v3_06 idlspec2d (sos3) --- v5_4_12 plate-mapper3 ------ NOCVS:v5_0_7 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5100 -5.5100 Axial B 16.0940 16.0940 Axial C 15.9210 15.9220 Trans D 11.0090 11.0100 Lateral E 10.6553 10.6553 Lateral F 12.2301 12.2301 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13053. 12647. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 18.4 Alt.: 61.765420 MIGS TONIGHT NOMINAL Axial A -8.4310 -7.8320 Axial B -0.6390 -0.0410 Axial C 13.2210 13.8180 Trans D 14.1930 14.1840 GALILS Commanded: 2224573. 2205846. 2238142. 3900. 5200. Actual: 2224522. 2205867. 2238052. 3882. 5210. SETMIR VALUES SecDesOrient: 1764.28 -9.00 -4.00 -58.00 -100.12 SecOrient: 1764.31 -9.03 -4.02 -57.69 -100.19